OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 14 -3x + 2y = 1

OpenStudy (dan815):

if you add both equations the x will dissappear

OpenStudy (dan815):

3x + y = 14 -3x + 2y = 1 -------------- 0X + 3y = 15 3y=15 y=15/3=5

OpenStudy (anonymous):

it cant be those though..

OpenStudy (anonymous):

5 3 y can be any number as there are infinitely many solutions to the system. There is no y value as there is no solution to the system.

OpenStudy (anonymous):

hello?